Yolo and Solano are two countries that both produce shrimp and tomatoes. In Yolo each worker in a one-hour period can produce either 28.95 pounds of shrimp or 12.96 pounds of vegetables. In Solano each worker in a one-hour period can produce either 27.72 pounds of shrimp or 4.93 pounds of tomatoes. Suppose both countries have constant opportunity cost of production and decide to specialize and exchange. The country that specializes in tomatoes is willing to sell 100 pounds of tomatoes for at least _____ pounds of shrimp.

Round to the nearest whole number
Answer #1

Opportunity cost of tomato in Yolo = 28.95/12.96 = 2.24 pounds of shrimp

Opportunity cost of tomato in Solano = 27.72/4.93 = 5.62 pounds of shrimp

Since the opportunity cost of tomato is lower for Yolo, it specializes in the production of tomatoes.

Thus, if it sells 100 tomatoes, it will need at least 2.24*100 = 224 shirmps